OpenStudy (anonymous):

Find the integral of sqrt{1-x^2}/x^2 dx ?

OpenStudy (anonymous):

\[\int\limits_{}^{} \sqrt{1-x^2}/x^2 dx\]

OpenStudy (astrophysics):

\[\int\limits \frac{ \sqrt{1-x^2} }{ x^2 }dx?\]

OpenStudy (astrophysics):

If so you can do a u sub \[u= \sqrt{1-x^2}\] and it seems fair to use by parts :)

OpenStudy (anonymous):

Okay I'll try that...

OpenStudy (anonymous):

Why would we not use trig substitution here?

OpenStudy (astrophysics):

You can try, I just thought this may be easier.

OpenStudy (anonymous):

Okay, can we try both ways? I want to get as much practice as possible :)

OpenStudy (astrophysics):

Sure, try both and see if they work, good way to check your answer s:)

OpenStudy (astrophysics):

\[\sqrt{a^2-x^2} \implies x = a \sin(t)\]

OpenStudy (astrophysics):

Did you try it out?

OpenStudy (anonymous):

Well I'm stuck on finding the integral of 1/sqrt{1+x^2}

OpenStudy (anonymous):

OH wait, I got it it's sin^(-1) (x)

OpenStudy (jhannybean):

\[\checkmark\]

OpenStudy (anonymous):

Wait so is it since 1/sinx is equal to sin^(-1) x

OpenStudy (astrophysics):

No, inverse does not = 1/x

OpenStudy (astrophysics):

\[f^{-1}(x) \neq \frac{ 1 }{ f(x) }\] haha I just showed this to the question below

OpenStudy (anonymous):

Okay, then I'm confused on how to get that...

OpenStudy (anonymous):

how to receive the integral of 1/sqrt{1-x^2}

OpenStudy (astrophysics):

You have to use by parts

OpenStudy (astrophysics):

\[u = \sqrt{1-x^2} ~~~du = -\frac{ x }{ \sqrt{1-x^2} }dx~~~,v = -\frac{ 1 }{ x } ~~~dv = \frac{ 1 }{ x^2 }dx\]

OpenStudy (astrophysics):

\[\int\limits u dv = uv - \int\limits v du\]

OpenStudy (astrophysics):

\[ \implies \sqrt{1-x^2}(-\frac{ 1 }{ x }) - \int\limits \left(- \frac{ x }{ \sqrt{1-x^2} } \right) (-\frac{ 1 }{ x })dx\]

OpenStudy (astrophysics):

\[ \implies -\frac{ \sqrt{1-x^2} }{ x } - \int\limits \frac{ 1 }{ \sqrt{1-x^2} }dx \implies -\frac{ \sqrt{1-x^2} }{ x }-\sin^{-1}(x)+C\]

OpenStudy (astrophysics):

Does that make sense, it may look confusing, just let me know if you don't understand something, try to redo it on paper as well.

OpenStudy (anonymous):

How did you find out that the integral of 1/sqrt{1-x^2} is the negative inverse of sin?

OpenStudy (astrophysics):

What is the derivative of arcsin?